OpenStudy (anonymous):

Simplify: 2 9/20+31/50

OpenStudy (anonymous):

Find a common multiple for 20 and 50. 100 works. So, 9/20 * 5/5 = 45/100 31/50 * 2/2 = 62/100 Add to get 107/100, or 1 7/100 If that was 2 9/20, then your final answer should be 3 7/100

OpenStudy (anonymous):

ok thanks